OpenStudy (fmnsmythe):

will fan and medal!!! What is the ∆G for the following reaction under standard conditions (T = 298 K) for the formation of NH4NO3(s)? 2NH3(g) + 2O2(g) NH4NO3(s) + H2O(l) Given: NH4NO3(s): ∆Hf = -365.56 kJ ∆Sf = 151.08 J/K. NH3(g): ∆Hf = -46.11 kJ ∆Sf = 192.45 J/K. H2O(l): ∆Hf = -285.830 kJ ∆Sf = 69.91 J/K. O2(g): ∆Hf = 0.00 kJ ∆Sf = 205 J/K. a. 186.6 kJ b. 6.9 kJ c. -10.4 kJ d. -126.3 kJ e. -382 kJ

OpenStudy (fmnsmythe):

Please help!

OpenStudy (amtran_bus):

Delta G = Delta H - T Delta S

OpenStudy (frostbite):

You could help a little by saying that he needs to calculate the enthalpy and entropy OF THE REACTION in order to calculate the Gibbs free energy OF THE REACTION.

OpenStudy (sweetburger):

\[\Delta H=\sum_{}^{}\Delta H_\left( products \right)-\sum_{}^{}\Delta H_\left( reactants \right)\] and \[\Delta S =\sum_{}^{}\Delta S_\left( products \right)-\sum_{}^{}\Delta S_\left( reactants \right)\]

OpenStudy (sweetburger):

these equations will be pertinent for this question